OpenStudy (anonymous):

Can anybody help me with a logarithmic scale plotting for given resistance and capacitance of a circuit ?

OpenStudy (anonymous):

Or maybe a link

OpenStudy (anonymous):

I find it a bit hard to see what your problem is since you did not really described your problem. If you have too little experience with logarithmic scales let Khan be your friend http://www.khanacademy.org/math/algebra/logarithms/v/logarithmic-scale If you need to know more about resistance and conductivity http://www.allaboutcircuits.com/vol_1/chpt_12/6.html Hope that helps.

OpenStudy (anonymous):

i have to plot a logarithmic scale for Uin/Uout on y axis and f/fo on x axis.fo=10^0,10^1.....until 10^6. f=1/2*pi*R*C

OpenStudy (anonymous):

i have problems on how to do calculation because at the cours were given some intermediate values 2 and 5 and i don't know if i use that values every time a want to plot a logarithmic scale.

OpenStudy (anonymous):

As the formula looks to me like something from a cut of frequency or low pass filter in electronics given that it should be c should be for capacity and not conductance, so I would start by checking that, so check if the intermediate values gives conductivity in some fraction of a Farad. http://en.wikipedia.org/wiki/Farad A bit more on the formula and more info on a electronic circuit I think we are dealing with. http://www.electronics-tutorials.ws/rc/rc_3.html Note: it is a long time since I had electronics in school, and I was not very motivated for school at the time, so I could be wrong.

OpenStudy (anonymous):

it's capacitance of course =)) i started this thing 2 weeks ago :D (my mistake)